# A Multi-agent LLM-based JUnit Test Generation with Strong Oracles

### v2

@@ Abstract
Unit testing plays a critical role in ensuring software correctness. However, writing unit tests manually is labor-intensive, especially for strong typed languages like Java, motivating the need for automated approaches. Traditional methods primarily rely on search-based or randomized algorithms to generate tests that achieve high code coverage and produce regression oracles, which are assertions derived from the program’s current behavior rather than its intended functionality. Recent advances in large language models (LLMs) have enabled oracle generation from natural language descriptions, aligning better with user requirements. However, existing LLM-based methods often require LLM fine-tuning or rely on external tools such as EvoSuite for test prefix generation.
@@
In this work, we propose CANDOR, a novel end-to-end, prompt engineering-based LLM framework for automated unit test generation in Java. CANDOR orchestrates multiple specialized LLM agents to collaboratively generate complete JUnit tests, including both high-quality test prefixes and accurate oracles. To mitigate the notorious hallucinations in LLMs and improve oracle correctness, we introduce a novel strategy that engages multiple reasoning LLMs in a panel discussion and generate accurate oracles based on consensus. Additionally, to reduce the verbosity of reasoning LLMs’ outputs, we propose a novel dual-LLM pipeline to produce concise and structured oracle evaluations.
@@
Our experiments on the HumanEvalJava and LeetCodeJava datasets show that CANDOR can generate accurate oracles and is slightly better than EvoSuite in generating tests with high line coverage and clearly superior in terms of mutation score. Moreover, CANDOR significantly outperforms the state-of-the-art, prompt-based test generator LLM-Empirical, achieving improvements of 15.8 to 25.1 percentage points in oracle correctness on both correct and faulty source code. Further ablation studies confirm the critical contributions of key agents in improving test prefix quality and oracle accuracy.
Keywords Automated Testing, Unit Test Generation, Large Language Models
1 Introduction
@@
Unit testing is a software testing activity wherein individual units of code are tested in isolation Siddiq et al. (2024). Unit testing plays a crucial role in modern software development because it helps software developers identify and fix defects in early phases Beck (2022). However, manually creating unit tests is laborious and necessitates substantial domain expertise to craft high-quality tests Daka et al. (2015); Daka and Fraser (2014). Consequently, developers often forgo writing tests for their code altogether. As reported in Siddiq et al. (2024); Gonzalez et al. (2017), only of studied Github projects contained test files.
To reduce the burden of manual test writing, various solutions have been proposed to automatically generate test suites Fraser and Arcuri (2011a); Pacheco et al. (2007); Shamshiri et al. (2015). A test suite typically comprises a set of test cases, where each test case is composed of a test prefix — inputs to the system under test (SUT) — and a test oracle that verifies whether the actual behavior matches the expected behavior of the SUT Hossain and Dwyer (2024). Prior works in automated test generation predominantly focuses on test prefix generation, with a goal of maximizing code coverage Tullis and Benjamin (2011); Serra et al. (2019). Techniques in this area include fuzzing Miller et al. (1990); Fioraldi et al. (2023), feedback-directed random test generation Csallner and Smaragdakis (2004); Pacheco and Ernst (2007); Pacheco et al. (2008); Selakovic et al. (2018); Arteca et al. (2022), dynamic symbolic execution Godefroid et al. (2005); Sen et al. (2005); Cadar et al. (2008); Tillmann et al. (2014), and search/evolutionary algorithm-based approaches Fraser and Arcuri (2011b, a). Although these techniques are successful in generating test suites with high code coverage, they are still not frequently employed in practice Fraser et al. (2015). Most of them rely on static or dynamic analysis to explore control and data flow paths in the program and consequently generate tests that maximize code coverage. One major limitation lies in their reliance on oracles that are automatically derived from the system’s current behavior. These are typical regression oracles, which assume the existing implementation is correct and simply record its behavior as the expected outcome. As such, they are primarily useful for detecting behavioral changes across software versions, not for validating functional correctness against intended specifications.
Recent advances in machine learning, especially the emergence of Large Language Models (LLMs), have inspired several works in generating test suites with specification-based oracles Hossain and Dwyer (2024); Dinella et al. (2022); Molina et al. (2024). Empirical studies have shown that LLMs can achieve state-of-the-art (SOTA) performance in oracle generation for weakly typed languages like JavaScript. For instance, a recent study Schäfer et al. (2023) found that LLMs can generate correct test oracles for JavaScript functions with high accuracy. However, test oracle generation for strongly typed languages such as Java remains significantly more challenging. This is due to the need for strict adherence to type constraints, language-specific syntax, and more complex execution semantics. To address this, recent work such as TOGLL Hossain and Dwyer (2024) combines EvoSuite-generated Fraser and Arcuri (2011a) test prefixes with a fine-tuned LLM for oracle generation. While effective, this hybrid approach has several limitations. First, it relies on EvoSuite to produce initial test scaffolds, which inherits the known limitations of EvoSuite, such as brittle test structures and inadequate test diversity Lemieux et al. (2023). Second, it requires fine-tuning on Java test data, which is costly to collect, often unavailable, and may not generalize well to new projects or domains.
In contrast, we advocate for a purely prompt-based approach using off-the-shelf LLMs, which is more practical, and widely accessible. Siddiq et al. Siddiq et al. (2024) explored this direction by using LLMs to perform end-to-end test generation for Java. Their method achieved lower line coverage on the HumanEvalJava benchmark than EvoSuite, and suffered from low oracle correctness. This is largely due to the uncertainty and hallucination of LLMs, which often produce assertions that appear plausible but are semantically incorrect.
To tackle these challenges, we propose a novel approach, CANDOR, for automated JUnit test generation with accurate oracles using off-the-shelf LLMs. CANDOR is an end-to-end test generation framework that eliminates the dependency on EvoSuite for test prefix generation, relies entirely on off-the-shelf LLMs without fine-tuning.
@@
For test prefix generation, CANDOR coordinates multiple specialized LLM agents—Initializer, Planner, Tester, and Inspector—to iteratively construct and refine test files. These agents are responsible for generating an initial test scaffold, designing test plans to improve coverage, producing executable test cases, and inspecting test case quality, respectively. At this stage, oracles are generated directly from the source code, which may be faulty in practice, and thus the resulting assertions—referred to as tentative oracles—may be incorrect.
@@
To ensure oracle correctness, CANDOR introduces a novel ensemble-based strategy inspired by David Hume’s quote, “Truth springs from arguments amongst friends,”, which underscores that constructive dialogue and diverse perspective lead to robust conclusions. In this spirit, CANDOR employs a panel discussion-style approach where multiple Panelist agents, powered by reasoning LLMs such as DeepSeek R1 Guo et al. (2025), independently evaluate tentative oracles against requirements derived from the natural language description of the SUT. Despite their effectiveness, such LLMs often suffer from verbose and redundant outputs, known as the “overthinking phenomenon” Sui et al. (2025). To address this, CANDOR introduce a dual-LLM pipeline, where a basic LLM (i.e., without reasoning capability) agent Interpreter extracts and formats key insights from each Panelist’s output. Finally, a Curator agent, also a basic LLM, aggregates these interpretations from the pipelines, determines necessary corrections, and generates accurate oracles through consensus. This design mitigates hallucination and uncertainty in test oracles by enabling multiple LLM agents to cross-validate tentative oracles and reach a consensus, ensuring that the final oracles align more closely with the intended behavior described in the specification.
@@
We evaluate CANDOR in terms of the quality of both test prefix and oracles on the benchmarking dataset HumanEvalJava Athiwaratkun et al. (2022) and a Leetcode-derived dataset from a popular Github repository 202 (Accessed: 2025a) that includes code with two complexity levels. Experimental results show that CANDOR is slightly better than EvoSuite in generating high-coverage test prefixes and clearly better in terms of mutation score, while producing accurate specification-based oracles. CANDOR also outperforms the SOTA end-to-end, LLM-based method LLM-Empirical by at least 15.8 percentage points in terms of oracle correctness on both correct and faulty source code. We also assess the individual contribution of the key agents in CANDOR. Experimental results show that removing Planner leads to substantial decrease in line coverage (0.099), branch coverage (0.130) and mutation score (0.111), respectively. Removing Requirement Engineer, and the panel discussion decreases oracle correctness by at least 0.005 and 0.158, respectively.
@@
1.
To the best of our knowledge, CANDOR is the first multi-agent LLM framework for end-to-end JUnit test generation, where specialized agents collaborate to iteratively generate, validate, and refine test cases, and do so without relying on external tools like EvoSuite or LLM fine-tuning.
2.
@@
4.
Experimental results on HumanEvalJava and LeetCode demonstrate that CANDOR compares favorably with EvoSuite in generating high quality test prefixes with slightly better line coverage and clearly higher mutation scores, while producing specification-based oracles with high accuracy, substantially outperforming a SOTA baseline by at least 15.8 percentage points.
2 Running Example
In this section, we present the running example in Figure 1, which is adapted from the HumanEvalJava dataset Chen et al. (2021). For each subject under test (SUT), we assume the availability of two relevant pieces of information: the source code and its natural language description.
Description. This part provides a natural language summary of the SUT’s main functionality. We assume such a description is available, as programmers typically provide such documentation explaining their code Siddiq et al. (2024). This description should include the input, the core implementation logic and the expected output. As depicted in Figure 1, the description specifies the SUT should take a list of integers as input, implement element-wise transformation and return the sum of the transformed list as output. For the th element , the transformation rule is as in Equation 1.
@@ 2 Running Example
This example represents a realistic and challenging test generation scenario, where the SUT is faulty and thus unreliable for oracle derivation. In real-world development, especially during early-stage or iterative coding, it is common for source code to contain latent bugs or incomplete logic Ettles et al. (2018). The goal of this paper is to generate high quality tests with accurate oracles, by combining both the source code and its natural language description, thus ensuring that the generated oracles are aligned with the intended functionality.
3 Methodology
This section demonstrates the workflow of CANDOR, comprising three steps: Initialization, Test Prefix Generation and Oracle Fixing. In the Initialization step, an initial test file is generated from the source code and iteratively refined by fixing syntactic errors found by a validation process, resulting in a syntactically correct test file . This file, typically containing only a small number of test cases (often ), serves as a template for the subsequent steps since it conforms to the syntax and conventions of Java and the testing framework JUnit 5. However, usually has low code coverage, as the focus at this stage is on syntactic correctness rather than test completeness. To enhance code coverage, the Test Prefix Generation step generates additional valid test prefixes with tentative oracles, producing an expanded test file . Notably, both Steps I and II rely solely on the source code for test generation. As a result, if the source code is faulty, the tentative oracles in , , and may also be incorrect. To address this issue, the Oracle Fixing step conducts a panel discussion among multiple LLM agents to revise and correct the oracles in , yielding the final test file .
3.1 Step I: Initialization
This step aims to generate a syntactically correct initial test file. As shown in the first box of Figure 2, it involves two main components: an LLM agent called the Initializer and a Validation process. Together, they form an iterative loop to ensure the generation of a valid initial test file .

### v6

# Hallucination to Consensus: Multi-Agent LLMs for End-to-End JUnit Test Generation

@@ Abstract.
Unit testing plays a critical role in ensuring software correctness. However, writing unit tests manually is labor-intensive, especially for strongly typed languages like Java, motivating the need for automated approaches. Traditional methods primarily rely on search-based or randomized algorithms to generate tests that achieve high code coverage and produce regression oracles, which are assertions derived from the program’s current behavior rather than its intended functionality. Recent advances in large language models (LLMs) have enabled oracle generation from natural language descriptions, aligning better with user requirements. However, existing LLM-based methods often require LLM fine-tuning or rely on external tools such as EvoSuite for test prefix generation, making them costly or cumbersome to apply in practice.
@@
In this work, we propose CANDOR, a novel end-to-end, prompt engineering-based LLM framework for automated unit test generation in Java. CANDOR orchestrates multiple specialized LLM agents to collaboratively generate complete JUnit tests, including both high-quality test prefixes and accurate oracles. To mitigate the notorious hallucinations in LLMs and improve oracle correctness, we introduce a novel strategy that engages multiple reasoning LLMs in a panel discussion and generates accurate oracles based on consensus. Additionally, to reduce the verbosity of reasoning LLMs’ outputs, we propose a novel dual-LLM pipeline to produce concise and structured oracle evaluations.
@@
Our experiments on the HumanEvalJava and more complex LeetCodeJava datasets show that CANDOR is comparable with EvoSuite in generating tests with high code coverage and clearly superior in terms of mutation score. Moreover, our prompt engineering-based approach CANDOR significantly outperforms the state-of-the-art, fine-tuning-based oracle generator TOGLL by at least 21.1 percentage points in oracle correctness on both correct and faulty source code. Further ablation studies confirm the critical contributions of key agents in improving test prefix quality and oracle accuracy.
1. Introduction
@@
Unit testing is a software testing activity wherein individual units of code are tested in isolation (Siddiq2024UsingStudy). Unit testing plays a crucial role in modern software development because it helps software developers identify and fix defects in early phases (Beck2022TestExample). However, manually creating unit tests is laborious and necessitates substantial domain expertise to craft high-quality tests (Daka2015ModelingTests; Daka2014AProblems). Consequently, developers often forgo writing tests for their code altogether. As reported in (Siddiq2024UsingStudy; Gonzalez2017AComprehension), only of studied Github projects contained test files.
To reduce the burden of manual test writing, various solutions have been proposed to automatically generate test suites (Fraser2011EvoSuite:Software; Pacheco2007Feedback-directedGeneration; Shamshiri2015Dot; xinyi; jiahui). A test suite typically comprises a set of test cases, where each test case is composed of a test prefix — inputs to the system under test (SUT) — and a test oracle that verifies whether the actual behavior matches the expected behavior of the SUT (Hossain2024TOGLL:LLMs).
Prior works in automated test generation predominantly focus on test prefix generation, with a goal of maximizing code coverage (Tullis2011OnLearning; Serra2019OnLater). Traditional techniques in this area include fuzzing (Miller1990AnUtilities; Fioraldi2023DissectingEvaluation), feedback-directed random test generation (Csallner2004JCrasher:Java; Pacheco2007Randoop:Java; Pacheco2008FindingTesting; Selakovic2018TestLanguages; Arteca2022Nessie:Callbacks), dynamic symbolic execution (Godefroid2005DART:Testing; Sen2005CUTE:C; Cadar2008EXE:Death; Tillmann2014TransferringDigger), and search/evolutionary algorithm-based approaches (Fraser2011EvolutionarySuites; Fraser2011EvoSuite:Software). Such methods are generally designed to maximize code coverage but often struggle to produce meaningful and human-readable test prefixes, primarily due to their limited understanding of the semantics of focal methods (zhang2025large). Therefore, recent research has shifted to leveraging large language models (LLMs) to produce more practical and meaningful test prefixes (codegeneratione). However, tang2024chatgpt and yang2024evaluation found that search-based approaches like EvoSuite (Fraser2011EvoSuite:Software) remain the best test prefix generator, substantially outperforming LLM-based approaches by approximately in terms of code coverage on benchmark datasets.
In contrast to the increasing attention on prefix generation, test oracle generation remains underexplored. Most existing works circumvent this challenge by generating regression oracles (Fraser2011EvolutionarySuites; Fraser2011EvoSuite:Software; Schafer2024AnGeneration; lemieux2023codamosa; jain2025testforge; 10.1145/3643769), which are derived from the system’s current behavior. Regression oracles assume the existing code implementation is correct and simply record its behavior as the expected outcome. As such, they are primarily useful for detecting behavioral changes across software versions, not for validating functional correctness against intended specifications.
A more rigorous and practical alternative to regression oracles is specification-based oracles, which are derived from specifications. However, generating correct specification-based oracles poses significant challenges, primarily because it necessitates rigorous reasoning about program semantics and expected behaviors. To that end, several works have proposed leveraging LLMs to generate oracles from specifications. Empirical studies have shown that LLMs can achieve state-of-the-art (SOTA) performance in oracle generation for weakly typed languages like JavaScript (Schafer2024AnGeneration) and Python (wang2024chat). However, generating test oracles for strongly typed languages such as Java remains significantly more challenging. This is due to the need for strict adherence to type constraints, language-specific syntax, and more complex execution semantics. Several works have attempted to address these challenges, using either fine-tuning (FT) or prompt engineering (PE) strategies (chatassert; Hossain2024TOGLL:LLMs; catlm; endres2024can; Siddiq2024UsingStudy). FT adapts LLMs to specific datasets while PE leverages off-the-shelf LLMs through carefully crafted prompts. FT-based approaches generally yield higher oracle correctness compared to prompt-engineering-based approaches. For instance, LLM-Empirical (Siddiq2024UsingStudy), as a representative state-of-the-art PE approach, demonstrates limited oracle correctness compared to FT-based approaches. In particular, TOGLL achieves the best oracle correctness by fine-tuning CodeParrot LLM on the SF110 dataset (Hossain2024TOGLL:LLMs). However, we identify three key limitations of TOGLL: (1) It relies on EvoSuite to produce initial test scaffolds, which inherits the known limitations of EvoSuite in generating human-readable test cases and reasoning about complex programs (lemieux2023codamosa); evosuite2014 reported side effects (e.g., “state pollution”) on of tests generated by EvoSuite (altmayer2025coverup). Moreover, EvoSuite has not been maintained since 2021, resulting in incompatibility issues with newly-developed projects (chen2024chatunitest). (2) It requires fine-tuning on version-specific Java data, which is costly to collect, often unavailable, and may not generalize well to new projects. Specifically, TOGLL is fine-tuned on test cases generated using EvoSuite v1.0.6, Java 8, and Ant Builder. Applying TOGLL to other environments (e.g., EvoSuite 1.2.0, Java 11, Maven Builder) would necessitate collecting new datasets and re-fine-tuning or domain adaptation techniques; otherwise, its performance would likely degrade substantially (domain). (3) The optimal configuration of TOGLL takes both Java docstrings and method signatures as input. Docstrings are informal natural language specifications that are ubiquitous in modern software (endres2024can), whereas method signatures are less common in most software code bases. As demonstrated in (Hossain2024TOGLL:LLMs), however, the absence of method signatures leads to a decrease in oracle correctness. Thus, reliance on method signatures makes TOGLL less robust when such information is incomplete or unavailable, which represents a practical limitation of the approach.
To address the afore-mentioned challenges in test prefix and oracle generation, we propose a novel approach, CANDOR, for automated JUnit test generation using off-the-shelf LLMs. We highlight that CANDOR is the first multi-agent-based framework for testing Java methods using only off-the-shelf LLMs, thereby laying the foundation for future research on Java test generation. In particular, CANDOR aims to generate test prefixes with at least comparable code coverage and mutation scores as EvoSuite. More importantly, it aims to derive accurate specification-based oracles directly from informal Java docstrings (hereafter referred to as descriptions), thereby compensating for the lack of formal specifications in most software systems.
@@
For test prefix generation, CANDOR coordinates multiple specialized LLM agents—Initializer, Planner, Tester, and Inspector—to construct and refine test files iteratively. These agents are responsible for generating an initial test scaffold, designing test plans to improve coverage, producing executable test cases, and inspecting the quality of those test cases. At this stage, oracles are generated directly from the source code, which may be faulty in practice, and thus the resulting assertions—referred to as tentative oracles—may be incorrect.
@@
To ensure oracle correctness, CANDOR introduces a novel ensemble-based strategy inspired by David Hume’s quote, “Truth springs from arguments amongst friends”, which underscores that constructive dialogue and diverse perspectives lead to robust conclusions. In this spirit, CANDOR employs a panel discussion-style approach in which multiple Panelist agents, powered by reasoning LLMs such as DeepSeek R1 (DeepSeek-AI2025DeepSeek-R1:Learning), independently evaluate tentative oracles against requirements derived from the natural-language description of the SUT. Noticeably, CANDOR requires only an informal description of the SUT’s function instead of a formal specification, following the practices in prior works (Hossain2024TOGLL:LLMs; chatassert; tang2024chatgpt). This can be attributed to recent findings suggesting that LLMs are robust to noisy natural-language instructions. However, effective application of LLMs is non-trivial. Despite their effectiveness, LLMs often suffer from verbose and redundant outputs, known as the “overthinking phenomenon” (Sui2025StopModels). To address this, CANDOR introduces a dual-LLM pipeline, where a basic LLM (i.e., without reasoning capability) agent Interpreter extracts and formats key insights from each Panelist’s output. Finally, a Curator agent, also a basic LLM, aggregates these interpretations from the pipelines, determines necessary corrections, and generates accurate oracles through consensus. This design mitigates hallucination and uncertainty in test oracles by enabling multiple LLM agents to cross-validate tentative oracles and reach a consensus, ensuring that the final oracles align more closely with the intended behavior described in the specification.
@@
We evaluate CANDOR in terms of the quality of both test prefixes and oracles on two benchmarking datasets: HumanEvalJava (Athiwaratkun2023Multi-lingualModels) and LeetCodeJava. HumanEvalJava is an extensively studied dataset in the domain of automated test/code generation (jiang2023impact; Siddiq2024UsingStudy; chatassert; yuan; kang2025explainable; li2024hybrid; endres2024can; tian2025fixing). To further assess CANDOR’s generalizability, we introduce a new dataset, LeetCodeJava, we constructed from the popular online judgement system LeetCode (2025LeetCodeDataset), which consists of solutions to programming problems with varying difficulty levels. Experimental results show that CANDOR is comparable with EvoSuite in generating high-coverage test prefixes and clearly better in terms of mutation score, while producing accurate specification-based oracles. CANDOR also outperforms the SOTA, fine-tuning-based approach TOGLL by at least 21.1 percentage points in terms of oracle correctness. We remark that this comparison is inherently unfavorable to CANDOR as it uses off-the-shelf LLMs, while TOGLL is fine-tuned with extra data. Fine-tuning adapts LLMs with domain-specific data, generally exhibiting higher effectiveness than prompt engineering in various tasks (shin2023prompt; shin2025prompt; paul2023enhancing; hua2025research). Nevertheless, because TOGLL is the state-of-the-art approach for Java oracle generation, we include this comparison for completeness. We also assess the individual contribution of the key agents in CANDOR. Experimental results show that removing Planner leads to a substantial decrease in line coverage (0.050), branch coverage (0.046), and mutation score (0.070), respectively. Removing the Requirement Engineer agent and the panel discussion decreases oracle correctness by at least 0.007 and 0.067, respectively.
@@
To the best of our knowledge, CANDOR is the first multi-agent LLM framework for end-to-end JUnit test generation, where specialized agents collaborate to generate, validate, and refine test cases iteratively. Unlike most existing approaches, CANDOR does not rely on expensive LLM fine-tuning or external tools like EvoSuite, which has not been maintained since 2021 and is difficult to adapt to new language features.
@@
Experimental results on two benchmarks, HumanEvalJava and LeetCode, demonstrate that CANDOR compares favorably with EvoSuite in generating high-quality test prefixes with comparable line/branch coverage and clearly higher mutation scores, while producing specification-based oracles with high accuracy, substantially outperforming a SOTA baseline by at least 21.1 percentage points.
The rest of this paper is organized as follows. Section 2 introduces a running example that facilitates the presentation of CANDOR in Section 3. Section 4 describes the experimental design, and Section 5 reports the corresponding results. Section 6 discusses key insights and threats to validity. Section 7 reviews related work, and Section 8 concludes the paper.
2. Running Example
In this section, we present the running example in Figure 1, which is adapted from the HumanEvalJava dataset (Chen2021EvaluatingCode). For each subject under test (SUT), we assume the availability of two relevant pieces of information: the source code and its natural language description.
Description. This part provides a natural language summary of the SUT’s main functionality. We assume such a description is available, as programmers typically provide such documentation explaining their code (Siddiq2024UsingStudy). This description should include the input, the core implementation logic, and the expected output. As depicted in Figure 1, the description specifies that the SUT should take a list of integers as input, implement element-wise transformation, and return the sum of the transformed list as output. For the th element , the transformation rule is as in Equation 1.
@@ 2. Running Example
This example represents a realistic and challenging test generation scenario, where the SUT is faulty and thus unreliable for deriving an oracle. In real-world development, especially during early-stage or iterative coding, it is common for source code to contain latent bugs or incomplete logic (Ettles2018CommonProgrammers). The goal of this paper is to generate high-quality tests with accurate oracles by combining both the source code and its natural language description, thus ensuring that the generated oracles are aligned with the intended functionality.
3. Methodology
This section demonstrates the workflow of CANDOR, comprising three steps: Initialization, Test Prefix Generation, and Oracle Fixing. In the Initialization step, an initial test file is generated from the source code and iteratively refined by fixing syntactic errors found by a validation process, resulting in a syntactically correct test file . This file, typically containing only a small number of test cases (often ), serves as a template for the subsequent steps since it conforms to the syntax and conventions of Java and the testing framework JUnit 5. However, usually has low code coverage, as the focus at this stage is on syntactic correctness rather than test completeness. To enhance code coverage, the Test Prefix Generation step generates additional valid test prefixes with tentative oracles, producing an expanded test file . Notably, both Steps I and II rely solely on the source code for test generation. As a result, if the source code is faulty, the tentative oracles in , , and may also be incorrect. To address this issue, the Oracle Fixing step conducts a panel discussion among multiple LLM agents to revise and correct the oracles in , yielding the final test file .
3.1. Step I: Initialization
This step aims to generate a syntactically correct initial test file. As shown in the first box of Figure 2, it involves two main components: (1) an LLM agent called the Initializer and (2) a Validation process. Together, they form an iterative loop to ensure the generation of a valid initial test file .